# Sparse Hanson-Wright Inequalities with Applications

### v2

@@ 5 Proofs of Theorem 3 and Corollary 1
With random vector $X$ defined in Theorem 3 and any positive even number $r$, we have the moment bound
$\mathbb{E}\left|\sum_{i=1}^{n}a_{i}X_{i}\right|^{r}$ $\leq(C_{\alpha}K^{2})^{r}\max\left\{r\|a\|_{\infty},\sqrt{r\sum_{i=1}^{n}a_{i}^{2}p_{i}}\right\}^{r}$
Proof.
Note that for a single variable we have $\mathbb{E}|X_{i}|^{l}\leq p_{l}K^{l}l^{\frac{l}{\alpha}}$. Let $\mathcal{P}(r,k)$ denote the set of all the partitions of $[r]$ into $k$ subsets and $\mathcal{P}_{\geq 2}(r,k)$ denote the subset of those partitions where all blocks have size at least 2. Furthermore, let $\lambda\coloneqq\sqrt{\sum_{i=1}^{n}a_{i}^{2}p_{i}}$. Since the $Y_{i}$’s are centered and satisfy the moment bounds in (5.1), we then obtain the following estimates:
@@
$\underset{(a)}{\leq}\sum_{k=1}^{r/2}\sum_{\Pi\in\mathcal{P}_{\geq 2}(r,k)}\prod_{\tau\in\Pi}\left(K^{|\tau|}|\tau|^{\frac{|\tau|}{\alpha}}\sum_{s=1}^{n}\|a\|_{\infty}^{|\tau|-2}a_{s}^{2}p_{s}\right)$ (5.5)
@@
$\underset{(b)}{\leq}K^{r}\sum_{k=1}^{r/2}\|a\|_{\infty}^{r-2k}\lambda^{2k}\frac{1}{k!}\sum_{\Pi\in\mathcal{P}_{\geq 2}(r,k)}\sum_{\phi\in\Phi(\Pi)}\left(\prod_{i=1}^{k}|\phi^{-1}(i)|^{|\phi^{-1}(i)|}\right)^{\frac{1}{\alpha}}$ (5.7)
=$\underset{(c)}{=}K^{r}\sum_{k=1}^{r/2}\|a\|_{\infty}^{r-2k}\lambda^{2k}\frac{1}{k!}\sum_{\begin{subarray}{c}\phi:[r]\to[k],\\ |\phi^{-1}(i)|\geq 2\end{subarray}}\left(\prod_{i=1}^{k}|\phi^{-1}(i)|^{|\phi^{-1}(i)|}\right)^{\frac{1}{\alpha}}$ (5.8)
@@
$\leq(C_{\alpha}K)^{r}\sum_{k=1}^{r/2}\|a\|_{\infty}^{r-2k}\lambda^{2k}\frac{r^{r}}{k^{k}}\sum_{r_{1}+\dots+r_{k}=r;r_{i}\geq 2}\left(\prod_{i=1}^{k}r_{i}^{r_{i}}\right)^{\frac{1}{\alpha}-1}.$ (5.10)
In the inequality (a) above, we use the the fact that $|\tau|\geq 2$ for $\tau\in\Pi$ so $a_{s}^{|\tau|-2}\leq\|a\|_{\infty}^{|\tau|-2}$. In the inequality (b), similar to the proof of Lemma 4.2, we use $\Phi(\Pi)$ to denote the set of surjections $\phi:[r]\to[k]$ corresponding to the partition $\Pi$, and further in (c) we rewrite the sum over $\Pi$ to surjections $\phi$.
Depending on the value of $\alpha$, we have the following two cases:
@@
    $\mathbb{E}\left|\sum_{i=1}^{n}a_{i}X_{i}\right|^{r}$ $\leq(C_{\alpha}K)^{r}\sum_{k=1}^{r/2}\|a\|_{\infty}^{r-2k}\lambda^{2k}\cdot\frac{r^{r}}{k^{k}}\sum_{r_{1}+\dots+r_{k}=r;r_{i}\geq 2}\left(\prod_{i=1}^{k}r_{i}^{r_{i}}\right)^{\frac{1}{\alpha}-1}$ (5.11)
@@
    $\leq(C_{\alpha}K)^{r}\sum_{k=1}^{r/2}\|a\|_{\infty}^{r-2k}\lambda^{2k}\cdot\frac{r^{r}}{k^{k}}\binom{r-1}{k-1}\left(4^{k}r^{r-2k+2}\right)^{\frac{1}{\alpha}-1}$ (5.12)
    $\leq(C_{\alpha}K^{2})^{r}\sum_{k=1}^{r/2}\|a\|_{\infty}^{r-2k}\lambda^{2k}\cdot r^{\frac{1}{\alpha}r-(\frac{2}{\alpha}+1)k}$ (5.13)
    $\leq(C_{\alpha}K)^{r}\sum_{k=1}^{r/2}(r^{\frac{1}{\alpha}}\|a\|_{\infty})^{r-2k}(\sqrt{r}\lambda)^{2k}$ (5.14)
@@
    $\underset{(d)}{\leq}(C_{\alpha}K)^{r}\max_{1\leq k\leq r/2}\max\left\{\frac{(r^{\frac{1}{\alpha}}\|a\|_{\infty})^{r}}{r/(r-2k)},\frac{(\sqrt{r}\lambda)^{r}}{r/(2k)}\right\}$ (5.15)
    $\leq(C_{\alpha}K)^{r}\max\left\{r^{\frac{1}{\alpha}}\|a\|_{\infty},\sqrt{r}\lambda\right\}^{r},$ (5.16)
    where the inequality (d) is due to Young’s inequality.
@@
  • When $\frac{1}{\alpha}<1$, simply bounding the term of $(\prod_{i=1}^{k}r_{i}^{r_{i}})^{\frac{1}{\alpha}-1}$ by 1 and using the Young’s inequality, we have
@@
    $\mathbb{E}\left|\sum_{i=1}^{n}a_{i}X_{i}\right|^{r}$ $\leq(C_{\alpha}K^{2})^{r}\sum_{k=1}^{r/2}\|a\|_{\infty}^{r-2k}\lambda^{2k}\cdot\frac{r^{r}}{k^{k}}\sum_{r_{1}+\dots+r_{k}=r;r_{i}\geq 2}\left(\prod_{i=1}^{k}r_{i}^{r_{i}}\right)^{\frac{1}{\alpha}-1}$
    $\leq(C_{\alpha}K^{2})^{r}\sum_{k=1}^{r/2}\|a\|_{\infty}^{r-2k}\lambda^{2k}\cdot{r^{r-k}}\binom{r-1}{k-1}$
@@
    $\leq(C_{\alpha}K^{2})^{r}\max\left\{r\|a\|_{\infty},\sqrt{r}\lambda\right\}^{r}$ (5.17)
Finally, Theorem 3 can be proved similarly to the last part of the proof of Theorem 1. With $C^{\prime}=e^{2}$, either the right-hand side of the inequality is at least one, or we can set $r$ to be the largest positive even integer satisfying

### v3

@@ 5.4 Proof of Theorem 2
Under the assumptions of Theorem 2, let $r$ be an even integer. We have
$\left\|\sum_{i=1}^{n}a_{i}X_{i}\right\|_{L_{r}}$ $\leq C_{\alpha}^{\prime}K\max\left\{r^{\max\{1,\frac{1}{\alpha}\}}\|a\|_{\infty},\sqrt{r}\lambda\right\}.$
Here, $C_{1}^{\prime}=8^{1+1/\alpha}e^{2}$.
Proof of Lemma 5.3.
Note that for a single variable we have $\mathbb{E}|X_{i}|^{l}\leq p_{i}K^{l}l^{\frac{l}{\alpha}}$. Let $\mathcal{P}(r,k)$ denote the set of all the partitions of $[r]$ into $k$ subsets and $\mathcal{P}_{\geq 2}(r,k)$ denote the subset of those partitions where all blocks have size at least 2. Recall that $\lambda\coloneqq\sqrt{\sum_{i=1}^{n}a_{i}^{2}p_{i}}$. We obtain the following estimates:
@@
$\underset{(\text{a})}{\leq}\sum_{k=1}^{r/2}\sum_{\Pi\in\mathcal{P}_{\geq 2}(r,k)}\prod_{\tau\in\Pi}\left(K^{|\tau|}|\tau|^{\frac{|\tau|}{\alpha}}\sum_{s=1}^{n}\|a\|_{\infty}^{|\tau|-2}a_{s}^{2}p_{s}\right)$ (5.52)
@@
$\underset{(\text{b})}{\leq}K^{r}\sum_{k=1}^{r/2}\|a\|_{\infty}^{r-2k}\lambda^{2k}\frac{1}{k!}\sum_{\Pi\in\mathcal{P}_{\geq 2}(r,k)}\sum_{\phi\in\Phi(\Pi)}\left(\prod_{i=1}^{k}|\phi^{-1}(i)|^{|\phi^{-1}(i)|}\right)^{\frac{1}{\alpha}}$ (5.54)
=$\underset{(\text{c})}{=}K^{r}\sum_{k=1}^{r/2}\|a\|_{\infty}^{r-2k}\lambda^{2k}\frac{1}{k!}\sum_{\begin{subarray}{c}\phi:[r]\to[k],\\ |\phi^{-1}(i)|\geq 2\end{subarray}}\left(\prod_{i=1}^{k}|\phi^{-1}(i)|^{|\phi^{-1}(i)|}\right)^{\frac{1}{\alpha}}$ (5.55)
@@
$\leq e^{2r}K^{r}\sum_{k=1}^{r/2}\|a\|_{\infty}^{r-2k}\lambda^{2k}\frac{r^{r}}{k^{k}}\sum_{r_{1}+\dots+r_{k}=r;r_{i}\geq 2}\left(\prod_{i=1}^{k}r_{i}^{r_{i}}\right)^{\frac{1}{\alpha}-1}.$ (5.57)
In the inequality (a) above, we use the the fact that $|\tau|\geq 2$ for $\tau\in\Pi$ so $a_{s}^{|\tau|-2}\leq\|a\|_{\infty}^{|\tau|-2}$. In the inequality (b), similar to the proof of Lemma 5.2, we use $\Phi(\Pi)$ to denote the set of surjections $\phi:[r]\to[k]$ corresponding to the partition $\Pi$, and further in (c) we rewrite the sum over $\Pi$ to surjections $\phi$. Depending on the value of $\alpha$, we have the following two cases:
@@
    $\mathbb{E}\left|\sum_{i=1}^{n}a_{i}X_{i}\right|^{r}$ $\leq e^{2r}K^{r}\sum_{k=1}^{r/2}\|a\|_{\infty}^{r-2k}\lambda^{2k}\cdot\frac{r^{r}}{k^{k}}\sum_{r_{1}+\dots+r_{k}=r;r_{i}\geq 2}\left(\prod_{i=1}^{k}r_{i}^{r_{i}}\right)^{\frac{1}{\alpha}-1}$ (5.58)
@@
    $\leq e^{2r}K^{r}\sum_{k=1}^{r/2}\|a\|_{\infty}^{r-2k}\lambda^{2k}\cdot\frac{r^{r}}{k^{k}}\binom{r-1}{k-1}\left(4^{k}r^{r-2k+2}\right)^{\frac{1}{\alpha}-1}$ (5.59)
    $\leq(2e^{2}K)^{r}\sum_{k=1}^{r/2}\|a\|_{\infty}^{r-2k}\lambda^{2k}\cdot\frac{r^{r}}{k^{k}}\left(4^{k}r^{r-2k+2}\right)^{\frac{1}{\alpha}-1}$ (5.60)
    $\underset{(\text{d})}{\leq}(2e^{2}K)^{r}(4r^{2})^{(\frac{1}{\alpha}-1)r}e^{\frac{r}{e}}\sum_{k=1}^{r/2}\|a\|_{\infty}^{r-2k}\lambda^{2k}\cdot r^{\frac{1}{\alpha}r-(\frac{2}{\alpha}+1)k}$ (5.61)
    $\leq(2e^{2+\frac{1}{e}}8^{\frac{1}{\alpha}-1}K)^{r}\sum_{k=1}^{r/2}(r^{\frac{1}{\alpha}}\|a\|_{\infty})^{r-2k}(\sqrt{r}\lambda)^{2k}$ (5.62)
@@
    $\underset{(\text{e})}{\leq}(2e^{2+\frac{1}{e}}8^{\frac{1}{\alpha}-1}K)^{r}\frac{r}{2}\cdot\max_{1\leq k\leq r/2}2^{r}\max\left\{\frac{(r^{\frac{1}{\alpha}}\|a\|_{\infty})^{r}}{r/(r-2k)},\frac{(\sqrt{r}\lambda)^{r}}{r/(2k)}\right\}$ (5.63)
    $\leq(8^{1/\alpha}e^{2}K)^{r}\max\left\{r^{\frac{1}{\alpha}}\|a\|_{\infty},\sqrt{r}\lambda\right\}^{r},$ (5.64)
    where the inequality (d) is by rearranging terms and apply $f(x)=(r/x)^{x}\leq e^{r/e}$ to the $k^{k}$ term and the inequality (e) is due to Young’s inequality.
@@
  • When $\frac{1}{\alpha}<1$, simply using the lower bound $(\prod_{i=1}^{k}r_{i}^{r_{i}})^{\frac{1}{\alpha}-1}\leq 1$ and the Young’s inequality, we have
@@
    $\mathbb{E}\left|\sum_{i=1}^{n}a_{i}X_{i}\right|^{r}$ $\leq(e^{2}K)^{r}\sum_{k=1}^{r/2}\|a\|_{\infty}^{r-2k}\lambda^{2k}\cdot\frac{r^{r}}{k^{k}}\sum_{r_{1}+\dots+r_{k}=r;r_{i}\geq 2}\left(\prod_{i=1}^{k}r_{i}^{r_{i}}\right)^{\frac{1}{\alpha}-1}$
    $\leq(e^{2}K)^{r}\sum_{k=1}^{r/2}\|a\|_{\infty}^{r-2k}\lambda^{2k}\cdot\frac{r^{r}}{k^{k}}\binom{r-1}{k-1}$
    $\leq(2e^{2}K)^{r}e^{r/e}\sum_{k=1}^{r/2}\|a\|_{\infty}^{r-2k}\lambda^{2k}\cdot r^{r-k}$
    $\leq(8e^{2}K)^{r}\max\left\{r\|a\|_{\infty},\sqrt{r}\lambda\right\}^{r}$ (5.65)
Therefore, combining the two cases, taking $C_{1}^{\prime}=8^{\max\{1,1/\alpha\}}e^{2}$ completes the proof. ∎
We are now positioned to prove Theorem 2. The proof follows a similar approach to that of Theorem 1 in Section 5.3. In essence, we define $r$ as the largest positive even integer satisfying
$2\leq r\leq\min\left\{\left(\frac{t}{C_{\alpha}^{\prime}eK\|a\|_{\infty}}\right)^{\min\{\alpha,1\}},\frac{t^{2}}{(C_{\alpha}^{\prime})^{2}e^{2}K^{2}\lambda^{2}}\right\}.$
By employing analogous arguments to those given in the proof of Theorem 1, we conclude the proof of Theorem 2 with $c_{\alpha}^{\prime}=(C_{\alpha}^{\prime}e)^{-2}$.
6 Proofs of Theorems 4 and 5
6.1 Proof of Theorem 4
We assume
$\sqrt{\sum_{i=1}^{n}a_{i}^{2}p_{i}}\leq\lambda\quad\text{and}\quad\|a\|_{\infty}\leq M.$
We begin by refining the moment estimates in Lemma 5.3 for $\alpha>1$. The existing moment bounds in Lemma 5.3 for $\alpha>1$ can be expressed as
@@ 6.1 Proof of Theorem 4
$\left\|\sum_{i=1}^{n}a_{i}X_{i}\right\|_{L_{r}}$$\leq\begin{cases}C_{\alpha}K(\sqrt{r}\lambda),&\text{if }r\leq\left(\frac{\lambda}{M}\right)^{2};\\ C_{\alpha}K(rM),&\text{if }r>\left(\frac{\lambda}{M}\right)^{2}.\end{cases}$ (6.1)
Our primary input is an improved moment estimate in the regime where $r>\left({\lambda}/{M}\right)^{2}$:
LEMMA 6.1.
Under the assumptions in Theorem 4, let $r$ be an even integer. We have